OpenStudy (anonymous):

very confused with this one. Someone please help! A package of modeling clay was shaped like a cone with a radius of 24 cm and a height of 6 cm. Miriam used all the clay to make a cylinder with a radius of 16 cm. What was the height of the cylinder? Use 3.14 to approximate pi, and express your final answer in tenths.

OpenStudy (amistre64):

determine volume of cone; and since volume remains constant, determine the parts ofr the volume of a cylindar and solve for h

OpenStudy (amistre64):

whats the formula for the volume of a cone?

OpenStudy (anonymous):

1/3 pi r^2 h

OpenStudy (amistre64):

good, and that of a cylindar?

OpenStudy (anonymous):

I don't know that one. Do you mind telling me so I can write it down?

OpenStudy (amistre64):

its just the cone without the 1/3: volume of cylindar is just, pi r^2h 1/3 pi (r1)^2 (h1) = pi (r2)^2 (h2); dividing by pi cancels out the common factor so i dont even know why you would need to approximate it 1/3 (r1)^2 (h1) = (r2)^2 (h2) ; now divide off the (r2)^2 1/3 (r1/r2)^2 (h1) = (h2)

OpenStudy (amistre64):

wht does that look like when we fill in the parts?